OpenStudy (anonymous):

WILL MEDAL AND FAN!! There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B. Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. (4 points) Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? (2 points) Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Jus

OpenStudy (prettygirl_shynice):

ill help

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i think this is part a PART A: a: 30*.20h=6 makes 36 houses b:45+3h= 48 houses

OpenStudy (anonymous):

@threshsupport

OpenStudy (threshsupport):

yes

OpenStudy (threshsupport):

A: mx=45(1.2)

OpenStudy (anonymous):

what does mx mean?

OpenStudy (tkhunny):

What? Try "A" again. 30 * 0.2= 6 -- You definitely do not have that.

OpenStudy (threshsupport):

ikr

OpenStudy (threshsupport):

i forgot how to set up the equation

OpenStudy (threshsupport):

been 2 years

OpenStudy (threshsupport):

wait shoot

OpenStudy (threshsupport):

i didn't read there were 2 neiborhoods

OpenStudy (tkhunny):

Try something with \((1.2)^{h}\).

OpenStudy (threshsupport):

neiborhood be would be mx=45+3 right?

OpenStudy (threshsupport):

ohhhh

OpenStudy (threshsupport):

man after you don't use the formulas for a while they go right out of your mind

OpenStudy (threshsupport):

mx=y

OpenStudy (anonymous):

sooo? i think this is part a PART A: a: 30*.20h=6 makes 36 houses b:45+3h= 48 houses

OpenStudy (threshsupport):

no

OpenStudy (tkhunny):

That's what you had before. Still no good.

OpenStudy (threshsupport):

the other person doesn't know who to set up the equation

OpenStudy (threshsupport):

wouldn't it be mx=45(1.2)^x and mx=45x+3

OpenStudy (threshsupport):

@tkhunny?

OpenStudy (tkhunny):

Closer. The first one starts with 30. The second should have the x on the 3.

OpenStudy (threshsupport):

so mx=30(1.2)^x and mx=45+3x

OpenStudy (anonymous):

? PART A: na:f(x) =30*0.2x nb:f(x)45+3x

OpenStudy (threshsupport):

part a is what i just said

OpenStudy (threshsupport):

the rest of your answers are right

OpenStudy (threshsupport):

for b

OpenStudy (anonymous):

part b? PART B: na:f(x)=30*0.2(5) =60 nb:f(x)=45+3(5) =60 the neighborhoods have the same number of houses after 5 years

OpenStudy (threshsupport):

yes

OpenStudy (threshsupport):

thats right :D gj

OpenStudy (anonymous):

PART C: after approximately 5 years the neighborhoods have the same amount of houses

OpenStudy (anonymous):

is that right too?

OpenStudy (anonymous):

@threshsupport

OpenStudy (tkhunny):

Still no. You haven't fixed it. f(x)=30*0.2(5) = 30*1 = 30 -- That is NOT what you want. \(f(x) = 30(1.2)^{x}\)-- This is a WHOLE different story.